OpenStudy (anonymous):

Could someone please help me understand this? Find the least common multiple of

OpenStudy (anonymous):

\[3x^{2}\] , \[12^{y}\] and \[10x^{3}y^{3}\]

OpenStudy (tester97):

when u said 12y was it supposed to be y to a power? or just y?

OpenStudy (anonymous):

woops sorry yeah its should be 12y

OpenStudy (tester97):

hmm if im remembering correctly would it be 3 for 3x^2 and 12y? and 2y for 12y and 10x^3y^3

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

so how'd you get that?

OpenStudy (e.mccormick):

A method to find them is to factor what is given. Then you mupliply back what is the union of them.

OpenStudy (anonymous):

okay, thanks guys!

OpenStudy (tester97):

no problem meh loe ^.^

OpenStudy (e.mccormick):

\(\color{red}{3\cdot x\cdot x}\) I need all 3 of these. \(\color{blue}{2\cdot 2\cdot} \color{red}{3}\color{blue}{\cdot y}\) I already have a 3. \(\color{green}{5}\cdot \color{blue}{2} \color{green}{\cdot x}\color{red}{\cdot x\cdot x} \color{green}{\cdot y\cdot y}\color{blue}{\cdot y}\) I already have a 2, 2 xes, an a y. \(\color{red}{3\cdot x\cdot x}\color{blue}{\cdot 2\cdot 2\cdot y}\color{green}{\cdot 5\cdot x\cdot y\cdot y}\) \(2\cdot 2\cdot 3\cdot 5\cdot x\cdot x\cdot x\cdot y\cdot y\cdot y\) \(60x^3y^3\)

OpenStudy (anonymous):

ohh okay thank you so much :D

OpenStudy (e.mccormick):

If I remember.... Let me see if I can fid a reference to it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thank you!!

OpenStudy (e.mccormick):

Their comment on highest power makes sense. Easier than what I did.

OpenStudy (e.mccormick):

My tri-color thing took a while to make. Hehe. So that is why the late detail. But I think it is worth it. Shows how I got the different parts of: \(\color{red}{3\cdot x\cdot x}\color{blue}{\cdot 2\cdot 2\cdot y}\color{green}{\cdot 5\cdot x\cdot y\cdot y}\)

OpenStudy (tester97):

so i just looked at that website and now that i think about it my answer is completely wrong sorry :(

OpenStudy (anonymous):

its okay Alley you tried :D And thank you @e.mccormick :D

OpenStudy (e.mccormick):

@tester97 Well, you got the start of it. You have to factor it to find the LCM. That is probably what you are remembering. How to start it. Now you can remember how to finish it too!

OpenStudy (tester97):

nahhhhhh mick i rather not cos i have something else to do right now i just came here cos katt said she needed help but i failed her........

OpenStudy (anonymous):

its okayyy Alley at least you tried! :D

OpenStudy (e.mccormick):

@tester97 /poke /poke Not that hard to remember this. Highest powers from each and then multiply out the union of the prime factors of the constant terms.

OpenStudy (tester97):

but but but mick......i dont wanna do anymore math right now im doing chemistry ugh...and btw i dont like to be poked :P

OpenStudy (e.mccormick):

Hmmm... well... union might not be 100% correct... eliminates some duplicates. Unique prime factors is more proper to say.

OpenStudy (e.mccormick):

Chemistry uses math!

OpenStudy (tester97):

lol i know but not too much

OpenStudy (e.mccormick):

Using math to blanace a chemical formula: http://youtu.be/yCxDAj87W8M http://saintjoe.edu/~karend/m244/ChemicalEquations.pdf

OpenStudy (tester97):

are u trying to give me a heart attack? im too pretty to die!

OpenStudy (e.mccormick):

Linear Algebra gets no love....